# Synthesizing Sentiment-Controlled Feedback For Multimodal Text and Image Data

### v3

@@ 4.2. Evaluation Metrics
Feedback synthesis is a one-to-many task, i.e., many feedbacks can be generated for one pair of image-text input. Hence, computing the accuracy of generated feedbacks is not feasible. Instead, we evaluate the generated feedbacks against the ground-truth comments using the metrics to evaluate semantic relevance and their ranks. For semantic relevance evaluation, Bilingual Evaluation Understudy (BLEU) (Papineni et al., 2002), Consensus-based Image Description Evaluation (CIDEr) (Vedantam et al., 2015), Recall-Oriented Understudy for Gisting Evaluation (ROUGE) (Lin, 2004), Semantic Propositional Image Captioning Evaluation (SPICE) (Anderson et al., 2016), and Metric for Evaluation of Translation with Explicit ORdering (METEOR) (Lavie and Denkowski, 2009) have been used. The BLEU is a precision-based metric, ROUGE is a recall-based metric, while METEOR combines precision and recall. On the other hand, CIDEr is a consensus-based evaluation metric, and SPICE is based on evaluating the sensitivity of n-grams. Higher values of these metrics denote more semantic similarity between the generated feedback and ground-truth comment. For ranking-based evaluation, we use ‘Mean Reciprocal Rank’ (Craswell, 2009) and ‘Recall@k’ (Runeson et al., 2007) defined below.
@@
Mean Reciprocal Rank: For calculating the Mean Reciprocal Rank (MRR), first, the similarity of the generated feedback is compared with all the ground-truth comments. If the most similar comment is ranked $k$, then the rank of the $j^{th}$ feedback is given by Eq. 10.
(10) $\begin{split}&rank_{j}=k\\ \end{split}$
where $k$ denotes the $k^{th}$ comment when sorted by the relevance whereas $rank_{j}$ is the Rank of the $j^{th}$ feedback. Finally, MRR is calculated as the average of the reciprocal ranks of all the generated feedback samples as per Eq. 11.
(11) $\begin{split}&MRR=(\frac{1}{n})\sum_{j=1}^{n}\frac{1}{rank_{j}}\\ \end{split}$
where $n$ is the number of generated feedback samples, while $rank_{j}$ denotes the $j^{th}$ feedback’s rank.
Recall@k: Recall@k counts the number of data samples matching any top-$k$ relevant data samples. Adapting Recall@k to evaluate the generated feedback, the number of feedbacks similar to any of the top-$k$ comments sorted according to relevance is calculated. To find if the generated feedback is similar to any comment, the rank of that feedback as calculated in Eq. 11 is used. Finally, the Recall@k can be formulated according to Eq. 12. According to this, if the rank of the feedback is in top-$k$, then a score of $1$ is assigned to the feedback, else 00. The summation of all scores is done to calculate the Recall@k as shown in Eq. 12.
(12) $\begin{split}&Recall@k_{i}=1\ \text{if}\ rank_{i}\in[1,\dots,k]\\ \vspace{.05in}&Recall@k=\sum_{i=1}^{n}Recall@k_{i}\end{split}$
@@
where $Recall@k_{i}$ and $rank_{i}$ show $i^{th}$ feedback’s Recall@k and rank, respectively and $Recall@k$ is the final score. Furthermore, the sentiments of the generated feedbacks have been computed and sentiment classification accuracy has been analysed along with the ‘Control Accuracy’ which is the difference between the accuracies of controlled and uncontrolled feedbacks.
4.3. Models
@@
The following constructed models have their architectures determined as per the ablation studies detailed in Section 4.4.7. Each includes the Controllability module outlined in Section 3.3.5, with the remaining architecture as follows.
@@
  • Baseline 3 implements a combination of a Transformer and a gated convolutional unit for textual encoding. It uses Faster RCNN with an additional visual attention mechanism for visual encoding. A late fusion strategy with averaging is used here.
  • Baseline 4 replaces the textual encoder with GPT-2 (OpenAI, 2019) and continues to Faster RCNN for visual data encoding with visual attention. It also combines the modalities using a late fusion with averaging. It has been empirically observed that GPT-$2$ based model generated good feedbacks only for textual input; however, it did not generate good feedbacks for multimodal input.
  • Proposed System incorporates Transformer as the textual encoder and Faster RCNN as the decoder and it uses concatenation along with late fusion.
4.4. Results
The results of the proposed benchmark feedback synthesis system are presented as follows.
4.4.1. Semantic Relevance Evaluation
The generated feedbacks’ semantic relevance with human comments has been evaluated. The feedbacks are generated to reflect the same sentiment class as reflected by the corresponding comments and then the feedbacks are evaluated using the BLEU, CIDEr, ROUGE, SPICE, and METEOR metrics. As depicted in Table 4, the proposed model has obtained the best values for these metrics in most cases.
Table 4. Semantic Relevance Evaluation.
@@ 4.4.1. Semantic Relevance Evaluation
4.4.2. Rank-based Evaluation
The generated feedbacks are evaluated using MRR and Recall@k. As observed in Table 5, $76.58$% feedbacks are relevant to one of the top $10$ comments and the MRR of $0.3789$ denotes that the generated feedbacks are contextually similar to one of the top $3$ comments.
[…]
@@ 4.4.2. Rank-based Evaluation
4.4.3. Sentiment-Control
Table 6 reports the Control Accuracies, which represent the difference in accuracies between controlled and uncontrolled feedbacks, for both the baselines and the proposed models. These models take the desired sentiment for the feedback to portray as one of the input parameters: 00 for negative and $1$ for positive. In uncontrolled settings, the parameter is not used and the control layer is disabled.
Table 6. Synthesized feedbacks’ sentiment analysis. Here, ‘USentiAcc’ and ‘CSentiAcc’ denote the sentiment classification accuracies for uncontrolled and controlled feedbacks respectively.
@@ 4.4.3. Sentiment-Control
The sentiment class of the feedback is determined using FLAIR (Akbik et al., 2019), SentimentR (Rinker, 2017), DistilBERT (Sanh et al., 2019), and RoBERTa (Liu et al., 2019), as detailed in Section 3.1. The model achieves a sentiment classification accuracy of $77.23\%$ and a control accuracy of $18.82\%$. To calculate the sentiment accuracy, one negative and one positive feedback is generated by passing the parameters—00 for negative and $1$ for positive. The sentiment of the feedback is then calculated and compared to the ground truth sentiment labels.
4.4.4. Compute Time Requirement
@@
The total amount of compute time required for one epoch of model training on V100100100100 GPU is as follows. Baseline $1$ required $168$ minutes, Baseline $2$ was the most time-consuming at $233$ minutes, Baseline $3$ needed $144$ minutes, and Baseline $4$ was the fastest at $103$ minutes. The Proposed System required $112$ minutes to complete one epoch.
4.4.5. Interpretability and Visualization
The model-level and case-level interpretability analyses have been incorporated. Model-level interpretability is achieved by introducing perturbations to the feedback synthesis model via the control layer. The impact of these perturbations on the output feedback is detailed in Table 6 in terms of sentiment classification accuracies for uncontrolled and controlled (perturbed) scenarios. Fig. 4 shows sample results, highlighting the features being focused on during the generation of uncontrolled and controlled feedbacks. In image plots, red and blue represent the most and least contributing pixels, respectively whereas for text plots, yellow and blue indicate the most and least important textual features. In Fig. 3(a), positive sentiments are indicated by smiling faces and a family setting, while negative sentiments are associated with the depiction of ageing, particularly in the older face. The expression of the girl on the left, a blend of smiling and discomfort, captures attention in both positive and negative contexts. The middle girl’s face, predominantly smiling, is highlighted in red for positive and blue for negative sentiments. Fig. 3(b) shows that dark areas contribute to negative sentiment, whereas faces are linked to positive sentiment. In negatively controlled settings, the crowd is focused; in positively controlled settings, the focus shifts to individual people. In Fig. 4(a), positive sentiments downplay the importance of the gun, concentrating more on the number plate. In the uncontrolled setting, the focus is primarily on the words. For Fig. 4(b), the facial features are highlighted red for positively controlled and blue for negatively controlled settings for the first image. The second image associates positive sentiment with light and text and negative sentiment with darkness.
(a) Sample Result 1
(b) Sample Result 2, continued on the next page.
(a) Sample Result 3
(b) Sample Result 4
Figure 4. Sample results along with interpretability plots. They depict the feedback generated by the proposed system using the news headline, text, and images (two out of multiple images shown) under the given sentiment-controllability constraint.
@@ 4.4.5. Interpretability and Visualization
Table 7. Human evaluation of generated feedbacks where $F_{UnCtrl}$, $F_{PosCtrl}$ and $F_{NegCtrl}$ show uncontrolled, positively and negatively controlled feedbacks. $Rel_{img}$, $Rel_{text}$, $Rel_{Comment}$ and $Rel_{F_{UnCtrl}}$ are ‘relevant with’ input images, text, comments and uncontrolled feedback, respectively.
@@
4.4.6. Human Evaluation
: The sentiments of the generated feedbacks have been evaluated by $50$ evaluators, comprising $25$ males and $25$ females, with an average age of $30\pm 2.73$ years. They assessed the controlled and uncontrolled feedbacks for their valence and relevance with the inputs. A total of 50 randomly picked samples have been evaluated, and the averages of the evaluators’ scores have been reported. These scores for the relevance ratings for three types of feedback—uncontrolled, positive, and negative—against images and texts have been described in Table 7. On average, $72.68$% and $78.14$% evaluators reported that the sentiments of positively and negatively controlled feedbacks are more positive and more negative respectively than uncontrolled feedbacks’. The higher relevance scores for controlled feedbacks ($F_{PosCtrl}$ and $F_{NegCtrl}$) compared to uncontrolled ones ($F_{UnCtrl}$) confirm the control layer’s influences in feedback’s desired sentiment alignment.
4.4.7. Ablation Studies
Following ablation studies have been conducted to evaluate the impact of various parameters on the proposed system’s performance.
1) Effect of Number of Control Layers and Value of Control-Parameter: The control layer has been used after the decoder and before the text generation phase to apply ‘control’ or constraints on the text generation. Here, it is crucial to decide a) the number of control layers and b) the suitable value for the control parameter. Regarding the number of control layers, we experimented with $1$, $2$, $3$, and $4$ control layers. The best performance has been observed using the $1$ control layer, which decreased slightly for $2$ control layers, decreased further for $3$ layers, and decreased significantly for $4$ layers. Further, regarding the control parameter, we experimented with its values of $5$%, $10$%, $15$%, and $20$%. The results show that more control can be achieved with the increasing value of the control parameter; however, fewer neurons will get trained for the entire training data, causing a degradation in the result’s quality. As depicted in Table 8, a control value of $10$% results in better performance than the other values. Hence, $1$ control layer with the control value of $10$% has been used in the final implementation.
2) Effect of Beam-size: The beam-size is a search parameter that refers to the number of options the model keeps at each step of the prediction, controlling the breadth of the search for the best output sequence. It keeps only the top $k$ predictions, where $k$ is the beam size. A larger beam size allows the model to explore more possibilities, potentially improving output quality; however, it increases computational requirements and may also degrade the output because of repetitive text generation. We experimented with beam-size values of $2$, $5$, $10$, $15$, and $20$. The corresponding sentiment classification accuracies and MRR values have been summarized in Table 8. As the beam-size value of $5$ provides the best performance and computational complexity trade-off, it has been used in the final implementation.
Table 8. Ablation studies on control parameter, x and beam-size. The entries show sentiment classification accuracy/MRR.
@@ 4.4.7. Ablation Studies
3) Effect of Division Factor for KAAP technique: The suitable values of the division factors $k_{img}$ and $k_{txt}$ used in Section 3.3.7 have been decided experimentally using the dice coefficient (Deng et al., 2018). It measures the similarity of two data samples; the value of $1$ denotes that the two compared data samples are completely similar, whereas a value of 00 denotes their complete dis-similarity. For each modality, we computed the KAAP values at $k\in\{2,3\dotsc,30\}$ and analyzed the dice coefficients for two adjacent $k$ values. For image & text, the dice coefficient values converge to $1$ at $k$ values of $5$ and $20$ respectively. Hence, the same have been used by the proposed system.
5. Discussion and Conclusion
In this section, we discuss the progression, draw conclusions from critical outcomes of our research, explore the limitations encountered, and outline the future directions for enhancing the controllability and effectiveness of feedback synthesis systems.
Advancements from Uncontrolled to Controllable Feedback Synthesis: The proposed controllable feedback synthesis system builds on our pioneering work on uncontrolled feedback synthesis (Kumar et al., 2023), which utilized a basic dataset collected from Twitter lacking detailed comment metadata. The current system enhances both the control and interpretability of feedback. It addresses the challenge of comment relevance in the CMFeed dataset by adopting Facebook’s relevance criteria. Despite slightly lower ‘R@k’ values due to more comments per sample ($10.5$ compared to $8.5$), the improved MRR value ($0.3789$ versus $0.3042$) highlights the effectiveness of the implemented similarity module. The proposed system, trained on human comments and image-text inputs under sentiment constraints, generates human-like feedback with appropriate sentiments, evidenced by metrics in Tables 4 and 5. The control layer facilitates the generation of feedback with desired sentiments, utilizing different non-keywords and varying keywords, especially at higher control parameter values.
Ethical Considerations, Data Privacy, and Potential Misuse: All data collection adheres to ethical guidelines with rigorous processes to ensure privacy and consent. Publicly available data has been manually scraped in accordance with Facebook’s ethical guidelines, which prohibit automatic scraping (Facebook, 2024b). Ethical protocols for scraping data from Facebook were strictly adhered to (Blog, 2024). Data from public news channels and Facebook posts is collected manually to comply with platform policies, and human-generated comments are included without identifying information. Furthermore, while any technology can be misused, we have proactively designed our feedback synthesis system with an integral interpretability module. This feature is crucial for explaining decision-making within the system and for detecting and preventing misuse, such as creating misleading content. Our goal is to ensure responsible and ethical use, enhancing positive impact while minimizing risks. By developing robust and transparent systems, we aim to foster trust and encourage responsible technology use in line with our ethical commitments.
Dataset Generalizability and Diversity: With $3.07$ billion of the total $5$ billion social media users, the extensive user base of Facebook, reflective of broader social media engagement patterns, ensures that the insights gained are applicable across various platforms, reducing bias and strengthening the generalizability of our findings. Additionally, the geographic and political diversity of these selected news sources, which range from local (NY Daily News) to international (BBC News) coverage and span the political spectrum from left (NY Daily News) to right (Fox News), ensures a balanced representation of global and political viewpoints in our dataset. This approach not only mitigates regional and ideological biases but also enriches the dataset with a wide array of perspectives, further solidifying the robustness and applicability of our research. Other large-scale news handles like NDTV, The Hindu, Xinhua, and SCMP were considered but not selected, as their content is often in regional languages such as Hindi or Chinese, which could limit the accessibility and relevance of the data.
Dataset Selection and Integration Considerations: We found a Reddit dataset containing human comments (Reddit, 2015), but opted not to include it in the considered datasets (Table 1) due to several issues. The source was deemed unreliable as it included a torrent download link and lacked associated scholarly publications. Moreover, the associated metadata was not comprehensive and did not appear to adhere to ethical protocols. These factors, along with the dataset’s unstructured and varied user-generated content and Reddit’s open forum format, which complicates data verification and includes potentially unreliable content, led to its exclusion. Furthermore, the informal nature of Reddit discussions does not align with the structured sentiment analysis required for news discourse. Additionally, we chose not to merge data from platforms such as Reddit and Twitter with Facebook due to the lack of comprehensive metadata, clear ethical guidelines, and control mechanisms—such as who can comment and whether users’ anonymity is maintained—on these platforms other than Facebook. These factors are critical for our analysis. Our focus on Facebook alone was crucial to ensure consistency in data quality and format.
Limitations: The efficacy of the CMFeed dataset is subject to several limitations, closely linked to data diversity and quality. Collection biases from specific platforms or demographics may impact generalizability. Technological constraints also limit the deployment of our system; substantial computational resources are required, which may not be readily available in all scenarios. Furthermore, adding modalities such as audio and physiological data may introduce privacy and ethical challenges. There is a risk of technology misuse in manipulating user sentiments and potential biases in sentiment analysis, which could result in over-reliance on automation, reduced human empathy, and unethical applications of feedback synthesis. These limitations highlight the need for cautious and responsible use of AI technologies, ensuring that ethical standards are rigorously maintained and biases are systematically addressed.

### v4

@@ IV-B Evaluation Metrics
Feedback synthesis is a one-to-many task, i.e., many feedbacks can be generated for one pair of image-text input. Hence, computing the accuracy of generated feedbacks is not feasible. Instead, we evaluate the generated feedbacks against the ground-truth comments using the metrics to evaluate semantic relevance and their ranks. For semantic relevance, we use BLEU [69], CIDEr [70], ROUGE [71], SPICE [72], and METEOR [73]. BLEU is precision-based, ROUGE is recall-based, METEOR combines both, CIDEr is consensus-based, and SPICE evaluates semantic structures or n-grams. Higher values of these metrics denote more semantic similarity between the feedback and ground-truth comment. For ranking-based evaluation, we use ‘Mean Reciprocal Rank’ and ‘Recall@k’.
@@
Mean Reciprocal Rank [74]: For Mean Reciprocal Rank (MRR), each generated feedback is compared with all ground-truth comments; if the most similar comment is ranked $k$, the $j^{th}$ feedback has $rank_{j}=k$ (the $k^{th}$ comment when sorted by relevance). MRR is the average reciprocal rank over all samples: $MRR=\frac{1}{n}\sum_{j=1}^{n}\frac{1}{rank_{j}}$, where $n$ is the number of generated feedback samples and $rank_{j}$ is the rank of the $j^{th}$ feedback.
@@
Recall@k [75]: Recall@k counts the number of data samples matching any top-$k$ relevant samples. For generated feedback, we count feedbacks similar to any of the top-$k$ comments sorted by relevance. If a feedback’s rank is in the top-$k$, it receives a score of $1$, else 0. Recall@k is then $\sum_{i=1}^{n}Recall@k_{i}$, where $Recall@k_{i}$ and $rank_{i}$ are the $i^{th}$ feedback’s Recall@k and rank. Sentiments of generated feedbacks are also computed, and sentiment classification accuracy is analyzed along with the ‘Control Accuracy’, the difference between accuracies of controlled and uncontrolled feedbacks.
IV-C Models
@@
The following models’ architectures are determined through ablation studies (Section IV-D7). Each includes the Controllability module, with the remaining architecture as follows
@@
  • Baseline 3 implements a combination of a Transformer and a gated convolutional unit for textual encoding. It uses Faster RCNN with visual attention mechanism for visual encoding and a late fusion strategy with averaging.
  • Baseline 4 replaces the textual encoder with GPT-2 [76] and continues to use Faster RCNN with visual attention for visual data encoding and late fusion with averaging. It has been empirically observed that GPT-$2$ based model generated good feedbacks only for textual input; however, it did not generate good feedbacks for multimodal input.
  • Proposed System incorporates Transformer as the textual encoder and Faster RCNN as visual encoder and it uses concatenation along with late fusion.
IV-D Results
IV-D1 Semantic Relevance Evaluation
The generated feedbacks’ semantic relevance with human comments has been evaluated. The feedbacks are generated to reflect the same sentiment class as reflected by the corresponding comments and then the feedbacks are evaluated using the BLEU, CIDEr, ROUGE, SPICE, and METEOR metrics. As depicted in Table IV, the proposed model has obtained the best values for these metrics in most cases.
TABLE IV: Semantic Relevance Evaluation.
@@ IV-D1 Semantic Relevance Evaluation
IV-D2 Rank-based Evaluation
The generated feedbacks are evaluated using MRR and Recall@k. As observed in Table V, 76.5876.58% feedbacks are relevant to one of the top $10$ comments and the MRR of $0.3789$ denotes that the generated feedbacks are contextually similar to one of the top $3$ comments.
TABLE V: Rank-based Evaluation. ‘MRR’ and ‘R@k’ denote ‘Mean Reciprocal Rank’ and ‘Recall@k’ where k $\in$ {1,3,5,10}.
@@ IV-D2 Rank-based Evaluation
IV-D3 Sentiment-Control
Table VI reports the Control Accuracies, which represent the difference in accuracies between controlled and uncontrolled feedbacks. In controlled settings, the desired sentiment for the feedback to portray (0 for negative and $1$ for positive) is one of the input parameters. In uncontrolled settings, the parameter is not used and the control layer is disabled.
TABLE VI: Synthesized feedbacks’ sentiment analysis. Here, ‘USentiAcc’ and ‘CSentiAcc’ denote the sentiment classification accuracies for uncontrolled and controlled feedbacks.
@@ IV-D3 Sentiment-Control
The sentiment class of the feedback is determined using FLAIR [59], SentimentR [60], DistilBERT [61], and RoBERTa [62], as detailed in Section III-A. The model achieves a sentiment classification accuracy of $77.23\%$ and a control accuracy of $18.82\%$. To calculate the sentiment accuracy, one negative and one positive feedback are generated by passing the parameters 0 for negative and $1$ for positive. The sentiment of the feedback is then calculated and compared to the ground truth sentiment labels.
IV-D4 Compute Time Requirement
@@
The total amount of compute time required for one epoch of model training on V100100 GPU is as follows. Baseline $1$ required $168$ minutes, Baseline $2$ was the most time-consuming at $233$ minutes, Baseline $3$ needed $144$ minutes, and Baseline $4$ was the fastest at $103$ minutes. The Proposed System required $112$ minutes to complete one epoch.
IV-D5 Interpretability and Visualization
The model-level and case-level interpretability analyses have been incorporated. Model-level interpretability is achieved by introducing perturbations to the feedback synthesis model via the control layer. The impact of these perturbations on the output feedback is detailed in Table VI in terms of sentiment classification accuracies for uncontrolled and controlled (perturbed) scenarios. Fig. 3 shows sample results with uncontrolled and controlled feedbacks while Supplementary Sec. S1 details pixel-level image attributions and word-level textual feature attributions.
Figure 3: Sample feedbacks generated by the proposed system using input text and images (one out of multiple images shown) with sentiment-control. Supplementary material’s Fig. S1 depicts feature heatmaps, salient words and color denotations.
IV-D6 Human Evaluation
The sentiments of the generated feedbacks were evaluated by $50$ evaluators ($25$ males, $25$ females; average age $30\pm 2.73$ years). They assessed controlled and uncontrolled feedbacks for valence and relevance. A total of $50$ randomly picked samples were evaluated, and averages of evaluators’ scores were reported (Table VII). On average, 72.6872.68% and 78.1478.14% reported that positively and negatively controlled feedbacks are more positive and more negative, respectively, than uncontrolled feedbacks. Higher relevance scores for controlled feedbacks ($F_{PosCtrl}$, $F_{NegCtrl}$) than uncontrolled ($F_{UnCtrl}$) confirm the control layer’s influence on desired sentiment alignment.
@@ IV-D6 Human Evaluation
TABLE VII: Human evaluation of generated feedbacks where $F_{UnCtrl}$, $F_{PosCtrl}$ and $F_{NegCtrl}$ show uncontrolled, positively and negatively controlled feedbacks. $Rel_{img}$, $Rel_{text}$, $Rel_{Comment}$ and $Rel_{F_{UnCtrl}}$ are ‘relevant with’ input images, text, comments and uncontrolled feedback, respectively.
@@
IV-D7 Ablation Studies
Following ablation studies evaluate various parameters’ impact on the proposed system’s performance.
Effect of Number of Control Layers and Value of Control-Parameter
The control layer is used after the decoder and before text generation to apply ‘control’ on text generation. It is crucial to decide (i) the number of control layers and (ii) the value of the control parameter. We experimented with $1$, 22, 33, and $4$ control layers. The best performance was with $1$ layer, decreasing slightly with $2$, further with $3$, and significantly with $4$. For the control parameter, we tried 55%, 1010%, 1515%, and 2020%. Increasing the value yields stronger control but trains fewer neurons on the full data, degrading quality. As shown in Table VIII, 1010% gives the best trade-off. Hence, we use $1$ control layer with a control value of 1010% in the final implementation.
Effect of Beam-size
The beam size is a search parameter referring to the number of options the model keeps at each prediction step, controlling the breadth of the search for the best output. It keeps only the top $k$ predictions, where $k$ is the beam size. A larger beam size allows exploration of more possibilities, improving output quality, but increasing computation and causing repetitive text generation. We experimented with beam sizes $2$, 55, $10$, $15$, and $20$; the corresponding sentiment classification accuracies and MRR values are shown in Table VIII. Beam size $5$ provided the best performance–complexity balance and is used in the final implementation.
TABLE VIII: Ablation studies on control parameter, x and beam-size. The entries show sentiment classification accuracy/MRR.
@@ IV-D7 Ablation Studies
The suitable values of the division factors $k_{img}$ and $k_{txt}$ used in Section III-C7 have been decided experimentally using the dice coefficient [77]. It measures two data samples’ similarity; the value of $1$ denotes that the two compared data samples are completely similar, whereas a value of 0 denotes their complete dis-similarity. For each modality, we computed the KAAP values at $k\in\{2,3\dotsc,30\}$ and analyzed the dice coefficients for two adjacent $k$ values. For image & text, the dice coefficient values converge to $1$ at $k$ of $5$ and $20$ respectively. Hence, the same have been used by the proposed system.
V Discussion and Conclusion
The proposed work’s contributions, outcomes, limitations, and future directions for controllable multimodal feedback synthesis are summarized as follows.
Advancements from Uncontrolled to Controllable Feedback Synthesis: Building on our prior work on uncontrolled feedback synthesis [1] utilizing basic data from Twitter without the comments metadata, we add sentiment control and interpretability. Using Facebook’s relevance criteria and richer comment metadata, CMFeed improves ranking (MRR from $0.3042$ to $0.3789$ under comparable settings) and, trained on human comments plus multimodal inputs, generates sentiment-aligned feedback (Tables IV, V). A control layer enables sentiment switching via targeted neuron activation.
Ethical Considerations, Data Privacy, and Potential Misuse: All data collection follows Facebook’s policies and established ethical guidelines [78, 79]. The process excludes personal identifiers and respects platform terms. The integrated interpretability module supports transparency and can help detect misuse, which reduces risks such as deceptive content and promotes responsible deployment.
Generalizability and Diversity: Facebook’s user base of $3.07$ billion provides broad coverage and demographic diversity. The four selected news sources (BBC News, Sky News, Fox News, and NY Daily News) span different regions and political orientations. This choice reduces bias and enriches the range of perspectives represented in the dataset.
Dataset Selection and Integration: We reviewed a Reddit dataset [80] but did not include it because of unreliable sourcing, incomplete metadata, and weak ethical safeguards. We also chose not to merge Reddit or Twitter with Facebook to preserve consistent metadata, clear usage policies, and privacy controls.
Limitations: The dataset may still contain platform-specific demographic biases. Training and deployment require significant computational resources. Adding audio or physiological modalities introduces additional privacy and ethics considerations. There is a risk of misuse for sentiment manipulation, which highlights the need for strict oversight and safeguards.